# Practical trainable temporal post-processor for multi-state quantum measurement

### v2

@@ V TPP learning: Simulation results
Crucially, when learning from data obtained from quantum systems, the observed correlations can have a quantum-mechanical origin. In what follows, we demonstrate the ability of the TPP to learn these quantum correlations, using simulations of two experimental setups where such quantum noise sources arise naturally: (i) readout using phase-preserving quantum amplifiers with a finite bandwidth, so that the amplifier added noise (demanded by quantum mechanics) has a nonzero correlation time, and (ii) readout of finite lifetime qubits with multi-level transitions (quantum jumps).
@@ V.2 Correlated quantum noise due to multi-level transitions
From readout datasets with no transitions to readout data with increasing transition rates, we note a small but clear improvement in classification performance using the trained TPP in comparison to the FGDA. That the TPP is able to learn information in the presence of transitions that evades the MF is clear when we compare the two sets of filters in Fig. 6(c). As the transition rates increase, the MF undergoes modifications due to the changes to the means of heterodyne records. However, the TPP is sensitive to changed beyond means - in the correlations of measured data - and increasingly learns a distinct filter with sharply decaying features. We note that the utility of similar exponential linear filters for finite-lifetime qubits has been the subject of earlier analytic work [21]. The TPP approach generalizes the ability to learn such filters in the presence of arbitrary transition rates and measurement tones, and for multi-state classification.
@@ VI Discussion and Outlook
The TPP approach, anchored by its connection to standard matched filtering under simplified readout conditions, with demonstrated advantages for real qubit readout under more complex readout conditions, and feasibility for FPGA implementations (to be demonstrated in future work), is ideal for integration with cQED measurement chains for the next step in readout optimization. Furthermore, the TPP’s generality and ability to learn from data could pave the way for an even broader class of applications. An important potential use is as a post-processor of quantum measurement data for quantum machine learning. With the use of general quantum machines for information processing, the optimal means to extract data from their measurements may not always be known. The TPP is ideally suited to uncover the optimal linear post-processing step, through training that could be incorporated parallel to, or as part of, the optimization of the quantum machine. Finally, optimal state estimation is essential for control applications. The trainable TPP can form part of a framework for control applications, such as Kalman filtering for quantum systems.
@@ Appendix A Experimental Setup
In this appendix section, we show a few more examples of readout IQ histograms as well as a more detailed circuit diagram for the measurement chain. Shown in Fig. 7 below, we see two examples of the extremes of the measurement data readout of Qubit B used to generate Fig. 3. Part (a) shows a lower power readout pulse performed for a short 300ns time, where the cavity barely has time to reach a steady state before the drive is turned off. Consequently, information from both the ring up and ring down must be integrated to achieve the SNR shown in this figure. Despite this measure, there is still significant infidelity from the lack of separation of the gaussian signals. In the second case, the displacement voltage is larger, and the pulse is three times as long, resulting in significantly increased separation of the gaussian signals and enabling discrimination of the |$\ket{g},\ket{e},\ket{f}$ and |$\ket{h}$ states. However, the large powers required induce transitions between these states, resulting in the trails between them as the measurement integrates a mixture of different cavity states at different times.
@@ B.1 Dispersive readout with no qubit transitions and using a cavity
The resulting SME is purely linear and can be solved exactly using a truncated equations of motion (TEOMs) approach.
@@ B.2 Dispersive readout with no qubit transitions and using a quantum-limited amplifier with added noise
As in the previous subsection, in the absence of any state transitions, $\mathcal{L}_{\rm envt}\to 0$, and the SME of Eq. (19) takes the simpler form:
@@
We now summarize the actual parameter choices used to generate quantum amplifier simulated datasets in the main text. We define the total cavity loss rate $\kappa=\kappa^{\prime}+\Gamma_{c}$. Then, we choose cavity parameters so that $\kappa^{\prime}=\Gamma_{c}=0.5\kappa$, and the dispersive shift $\chi/\kappa=0.5$. Recall that perfect non-reciprocal coupling in the desired direction requires $g=\Gamma=0.5\kappa$. Lastly, amplifier parameters are chosen so that $\gamma=\gamma_{d}+\Gamma=5\kappa$, yielding the ratio of cold amplifier to cavity linewidth $\gamma/\kappa=5$ used in the main text, and also implying that $\gamma_{d}=4.5\kappa$.
@@
for $\mathcal{L}_{\rm meas}$ now given by Eq. (35). The resulting SME is again linear and can be solved exactly using a truncated equations of motion (TEOMs) approach.
@@ B.3 Dispersive readout including multi-level transitions using a cavity
We emphasize that now the quantum state of the measurement chain can not generally be expressed in the form of Eq. (26). Hence Eq. (41) is integrated in joint the qubit-cavity Hilbert to generated simulated measurement datasets.
@@ C.1 Cost function and learned weights
where $\mathbf{I}$ is the identity matrix on $\mathbb{R}^{(N_{\rm O}\cdot N_{\rm T}+1)\times(N_{\rm O}\cdot N_{\rm T}+1)}$ and $\lambda$ is typically referred to as a regularization parameter. If $\mathbf{C}$ is invertible, we have $\mathbf{C}^{+}\to\mathbf{C}^{-1}$. We do emphasize that for the datasets analyzed in this paper, we consider a weak regularization of $\lambda=10^{-15}$, if any.
@@ C.2 Testing via cross-validation
For all classification infidelities calculated in the main text, we perform cross-validation. For a full dataset of $N_{\rm traj}$ records per state, a training set is constructed with $N_{\rm train}<N_{\rm traj}$ records as described above. The remaining $N_{\rm test}=N_{\rm traj}-N_{\rm train}$ records are used to construct a testing set. We use 80% of the dataset for training, and the remaining 20% for testing. This is consistent with training and testing set sizes for standard machine learning applications; for example, the MNIST handwritten digits classification task [61] uses 85.7% of the total dataset for training, and the remaining 14.3% for testing. Predicted state labels are obtained using this testing set via both the FGDA scheme, Eq. (10) of the main text, and the TPP, Eq. (1). This process is repeated a total of $L=10$ times: each time, a new set of weights $\bm{\mathcal{W}}^{\rm opt}$ is obtained from a distinct randomly chosen training set of the total $N_{\rm traj}$ records, and classification infidelities computed using the new random testing datasets. All classification fidelities are averaged to obtain the final values plotted in the main text. This cross-validation approach is standard in machine learning, and ensures that the observed performance is not unduly effected by variations due to the specific training or testing dataset used.
@@ C.3 Dependence on size and fidelity of training sets
Figure 9: Comparison of exact TPP versus FGDA as a function of training set size and qubit initialization error. Results are shown for training set sizes of (a) $N_{\rm train}=4000$ measurement records and (b) $N_{\rm train}=8000$ measurement records per class. More opaque markers indicate lower qubit initialization error.
@@ D.1.1 Simplification of mean matrix $\mathbf{M}$
$N_{\rm train}\mathbf{M}$ $=\mathbf{Y}\begin{pmatrix}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&\vec{1}^{T}\end{pmatrix}=\begin{pmatrix}\mathbf{Y}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&\mathbf{Y}\vec{1}^{T}\end{pmatrix}$ (64)
We now proceed to simplify the general matrices $\mathbf{Y}\vec{1}^{T}$ and $\mathbf{Y}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}$. Starting with the former, which simply yields a column vector that is an element of $\mathbb{R}^{C\times 1}$, we find explicitly:
@@
$\mathbf{M}=\begin{pmatrix}({\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(1)})^{T}&1\\ \vdots&\vdots\\ ({\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(C)})^{T}&1\end{pmatrix}\equiv\begin{pmatrix}(\vec{S}^{(1)})^{T}\\ \vdots\\ (\vec{S}^{(C)})^{T}\end{pmatrix},$ (67)
@@ D.1.2 Simplification of second-order moments matrix $\mathbf{C}$
$N_{\rm train}\mathbf{C}\equiv\bm{\mathcal{X}}\bm{\mathcal{X}}^{T}=\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\mathbf{X}\\ \vec{1}\end{pmatrix}\begin{pmatrix}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}^{T}&\vec{1}^{T}\end{pmatrix}=\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\mathbf{X}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}^{T}&{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\mathbf{X}\vec{1}^{T}\\ \vec{1}\mathbf{X}^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}^{T}&\vec{1}\vec{1}^{T}\end{pmatrix}$ (68)
@@
$\mathbf{C}=\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}(\mathbf{G}\!+\!\mathbf{V}){\color[rgb]{0.0,0.0,0.0}\mathbf{L}}^{T}&\sum_{c}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&C\end{pmatrix}$ (76)
@@
Having constructed explicit forms of $\mathbf{M}$ and $\mathbf{C}$, we are in principle positioned to evaluate the optimal weights $\bm{\mathcal{W}}^{\rm opt}$ explicitly as well. To do so, it first again proves useful to interpret the learned weights in terms of optimal filters.
@@ D.2 Constraints on TPP filters
$\bm{\mathcal{W}}^{\rm opt}\equiv\begin{pmatrix}(\bm{\vec{f}}_{1})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{-1}}&b_{1}&\\ \vdots&\vdots\\ (\bm{\vec{f}}_{C})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{-1}}&b_{C}\\ \end{pmatrix}\equiv\begin{pmatrix}(\vec{F}_{1})^{T}\\ \vdots\\ (\vec{F}_{C})^{T}\end{pmatrix}$ (77)
@@
$\mathbf{C}\vec{n}=\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}(\mathbf{G}\!+\!\mathbf{V}){\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&\sum_{c}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&C\end{pmatrix}\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}=\begin{pmatrix}\sum_{c}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ C\end{pmatrix}=\sum_{c}\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ 1\end{pmatrix}=\sum_{c}\vec{S}^{(c)}$ (81)
@@ D.3 Analytically-calculable TPP filters: “matched filters” for arbitrary $C$
$\mathbf{C}=\begin{pmatrix}\mathbf{L}\mathbf{G}\mathbf{L}^{T}\!+\!\mathbf{L}\mathbf{V}\mathbf{L}^{T}&\sum_{c}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&C\end{pmatrix}.$ (85)
@@
$\mathbf{C}=\begin{pmatrix}\mathbf{L}\mathbf{G}\mathbf{L}^{T}\!+\!\bar{\mathbf{I}}&\sum_{c}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&C\end{pmatrix}.$ (87)

### v3

@@ V TPP learning: Simulation results
Crucially, data obtained from quantum systems can exhibit temporal correlations that have a quantum-mechanical origin. In what follows, we demonstrate the ability of the TPP to learn these quantum correlations, using simulations of two experimental setups where such quantum noise sources arise naturally: (i) readout using phase-preserving quantum amplifiers with a finite bandwidth, so that the amplifier added noise (demanded by quantum mechanics) has a nonzero correlation time, and (ii) readout of finite lifetime qubits with multi-level transitions (quantum jumps).
@@ V.2 Correlated quantum noise due to multi-level transitions
From readout datasets with no transitions to readout data with increasing transition rates, we note a small but clear improvement in classification performance using the trained TPP in comparison to the FGDA. That the TPP is able to learn information in the presence of transitions that evades the MF is clear when we compare the two sets of filters in Fig. 6(c). As the transition rates increase, the MF undergoes modifications due to the changes to the means of heterodyne records. However, the TPP is sensitive to changes beyond means - in the correlations of measured data - and increasingly learns a distinct filter with sharply decaying features. We note that the utility of similar exponential linear filters for finite-lifetime qubits has been the subject of earlier analytic work [21]. The TPP approach generalizes the ability to learn such filters in the presence of arbitrary transition rates and measurement tones, and for multi-state classification.
@@ VI Discussion and Outlook
The TPP approach, anchored by its connection to standard matched filtering, with demonstrated advantages for real qubit readout under complex readout conditions, and feasibility for FPGA implementations (to be demonstrated in future work), is ideal for integration with cQED measurement chains for the next step in readout optimization. Furthermore, the TPP’s generality and ability to efficiently learn from data could pave the way for an even broader class of applications. An important potential use is as a post-processor of quantum measurement data for quantum machine learning. With the use of general quantum machines for information processing, the optimal means to extract data from their measurements may not always be known. We believe the TPP is ideally suited to uncover the optimal linear post-processing step, through training that could be incorporated as part of the optimization of the quantum machine. This is because the existence of an exact analytic form for the optimal trained TPP weights eliminates the need for multiple training epochs, batch-wise evaluations, or gradient computations, so that training the TPP adds minimal complexity to the optimization of an already complex quantum measurement chain, in stark contrast to the substantial overhead of training a neural network used as a post-processor. Finally, optimal state estimation is essential for control applications. The trainable TPP can form part of a framework for control applications, such as Kalman filtering for quantum systems.
@@ Appendix A Experimental Setup
In this appendix section, we show a few more examples of readout IQ histograms as well as a more detailed circuit diagram for the measurement chain. Shown in Fig. 7 below, we see two examples of the extremes of the measurement data for readout of Qubit B used to generate Fig. 3. Part (a) shows a lower power readout pulse performed for a short 300ns time, where the cavity barely has time to reach a steady state before the drive is turned off. Consequently, information from both the ring up and ring down must be integrated to achieve the SNR shown in this figure. Despite this measure, there is still significant infidelity from the lack of separation of the gaussian signals. In the second case, the displacement voltage is larger, and the pulse is three times as long, resulting in significantly increased separation of the gaussian signals and enabling discrimination of the |$\ket{g},\ket{e},\ket{f}$ and |$\ket{h}$ states. However, the large powers required induce transitions between these states, resulting in the trails between them as the measurement integrates a mixture of different cavity states at different times.
@@ B.1 Dispersive readout with no qubit transitions and using a cavity
The resulting SME preserves Gaussian states and can be solved exactly using a truncated equations of motion (TEOMs) approach.
@@ B.2 Dispersive readout with no qubit transitions and using a quantum-limited amplifier with added noise
In the absence of any state transitions, $\mathcal{L}_{\rm envt}\to 0$, and the SME of Eq. (19) takes the simpler form:
@@
We now summarize the actual parameter choices used to generate quantum amplifier simulated datasets in the main text. We define the total cavity loss rate $\kappa=\kappa^{\prime}+\Gamma$. Then, we choose cavity parameters so that $\kappa^{\prime}=\Gamma=0.5\kappa$, and the dispersive shift $\chi/\kappa=0.5$. Recall that perfect non-reciprocal coupling in the desired direction requires $g=\Gamma=0.5\kappa$. Lastly, amplifier parameters are chosen so that $\gamma=\gamma_{d}+\Gamma=5\kappa$, yielding the ratio of cold amplifier linewidth to cavity linewidth $\gamma/\kappa=5$ used in the main text, and also implying that $\gamma_{d}=4.5\kappa$.
@@
for $\mathcal{L}_{\rm meas}$ now given by Eq. (35). The resulting SME again preserves Gaussian states and can be solved exactly using a truncated equations of motion (TEOMs) approach.
@@ B.3 Dispersive readout including multi-level transitions using a cavity
We emphasize that now the quantum state of the measurement chain can not generally be expressed in the form of Eq. (26). Hence Eq. (41) is integrated in the joint qubit-cavity Hilbert space to generate simulated measurement datasets.
@@ C.1 Cost function and learned weights
where $\mathbf{I}$ is the identity matrix on $\mathbb{R}^{(N_{\rm O}\cdot N_{\rm T}+1)\times(N_{\rm O}\cdot N_{\rm T}+1)}$ and $\lambda$ is typically referred to as a regularization parameter. If $\mathbf{C}$ is invertible, we have $\mathbf{C}^{+}\to\mathbf{C}^{-1}$. We do emphasize that for the datasets analyzed in this paper, the intrinsic dataset noise serves as an effective regularizer, such that we can typically set $\lambda=0$.
@@ C.2 Testing via cross-validation
For all classification infidelities calculated in the main text, we perform cross-validation. For a full dataset of $N_{\rm traj}$ records per state, a training set is constructed with $N_{\rm train}<N_{\rm traj}$ records as described above. The remaining $N_{\rm test}=N_{\rm traj}-N_{\rm train}$ records are used to construct a testing set. We use 80% of the dataset for training, and the remaining 20% for testing. This is consistent with training and testing set sizes for standard machine learning applications; for example, the MNIST handwritten digits classification task [61] uses 85.7% of the total dataset for training, and the remaining 14.3% for testing. Predicted state labels are obtained using this testing set via both the FGDA scheme, Eq. (10) of the main text, and the TPP, Eq. (1). This process is repeated until a total of $L=10$ iterations are completed: each time, a new set of weights $\bm{\mathcal{W}}^{\rm opt}$ is obtained from a distinct randomly chosen training set of the total $N_{\rm traj}$ records, and classification infidelities computed using the new random testing datasets. All classification fidelities are averaged to obtain the final values plotted in the main text. This cross-validation approach is standard in machine learning, and ensures that the observed performance is not unduly effected by variations due to the specific training or testing dataset used.
@@ C.3 Dependence on size and fidelity of training sets
Figure 9: Comparison of general TPP versus FGDA as a function of training set size and qubit initialization error. Results are shown for training set sizes of (a) $N_{\rm train}=4000$ measurement records and (b) $N_{\rm train}=8000$ measurement records per class. More opaque markers indicate lower qubit initialization error.
@@ D.1.1 Simplification of mean matrix $\mathbf{M}$
$N_{\rm train}\mathbf{M}$ $=\mathbf{Y}\begin{pmatrix}\mathbf{X}^{T}\mathbf{L}^{T}&\vec{1}^{T}\end{pmatrix}=\begin{pmatrix}\mathbf{Y}\mathbf{X}^{T}\mathbf{L}^{T}&\mathbf{Y}\vec{1}^{T}\end{pmatrix}$ (64)
We now proceed to simplify the general matrices $\mathbf{Y}\vec{1}^{T}$ and $\mathbf{Y}\mathbf{X}^{T}\mathbf{L}^{T}$. Starting with the former, which simply yields a column vector that is an element of $\mathbb{R}^{C\times 1}$, we find explicitly:
@@
$\mathbf{M}=\begin{pmatrix}(\mathbf{L}\bm{\vec{s}}^{(1)})^{T}&1\\ \vdots&\vdots\\ (\mathbf{L}\bm{\vec{s}}^{(C)})^{T}&1\end{pmatrix}\equiv\begin{pmatrix}(\vec{S}^{(1)})^{T}\\ \vdots\\ (\vec{S}^{(C)})^{T}\end{pmatrix},$ (67)
@@ D.1.2 Simplification of second-order moments matrix $\mathbf{C}$
$N_{\rm train}\mathbf{C}\equiv\bm{\mathcal{X}}\bm{\mathcal{X}}^{T}=\begin{pmatrix}\mathbf{L}\mathbf{X}\\ \vec{1}\end{pmatrix}\begin{pmatrix}\mathbf{X}^{T}\mathbf{L}^{T}&\vec{1}^{T}\end{pmatrix}=\begin{pmatrix}\mathbf{L}\mathbf{X}\mathbf{X}^{T}\mathbf{L}^{T}&\mathbf{L}\mathbf{X}\vec{1}^{T}\\ \vec{1}\mathbf{X}^{T}\mathbf{L}^{T}&\vec{1}\vec{1}^{T}\end{pmatrix}$ (68)
@@
$\mathbf{C}=\begin{pmatrix}\mathbf{L}(\mathbf{G}\!+\!\mathbf{V})\mathbf{L}^{T}&\sum_{c}\mathbf{L}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}\mathbf{L}^{T}&C\end{pmatrix}$ (76)
@@
Having constructed explicit forms of $\mathbf{M}$ and $\mathbf{C}$, we are in principle positioned to evaluate the optimal weights and biases $\bm{\mathcal{W}}^{\rm opt}$ explicitly as well. To do so, it first again proves useful to interpret the learned weights in terms of optimal filters.
@@ D.2 Constraints on TPP filters
$\bm{\mathcal{W}}^{\rm opt}\equiv\begin{pmatrix}(\bm{\vec{f}}_{1})^{T}\mathbf{L}^{-1}&b_{1}&\\ \vdots&\vdots\\ (\bm{\vec{f}}_{C})^{T}\mathbf{L}^{-1}&b_{C}\\ \end{pmatrix}\equiv\begin{pmatrix}(\vec{F}_{1})^{T}\\ \vdots\\ (\vec{F}_{C})^{T}\end{pmatrix}$ (77)
@@
$\mathbf{C}\vec{n}=\begin{pmatrix}\mathbf{L}(\mathbf{G}\!+\!\mathbf{V})\mathbf{L}^{T}&\sum_{c}\mathbf{L}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}\mathbf{L}^{T}&C\end{pmatrix}\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}=\begin{pmatrix}\sum_{c}\mathbf{L}\bm{\vec{s}}^{(c)}\\ C\end{pmatrix}=\sum_{c}\begin{pmatrix}\mathbf{L}\bm{\vec{s}}^{(c)}\\ 1\end{pmatrix}=\sum_{c}\vec{S}^{(c)}$ (81)
@@ D.3 Analytically-calculable TPP filters: “matched filters” for arbitrary $C$
$\mathbf{C}=\begin{pmatrix}\mathbf{L}\mathbf{G}\mathbf{L}^{T}\!+\!\mathbf{L}\mathbf{V}\mathbf{L}^{T}&\sum_{c}\mathbf{L}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}\mathbf{L}^{T}&C\end{pmatrix}.$ (85)
@@
$\mathbf{C}=\begin{pmatrix}\mathbf{L}\mathbf{G}\mathbf{L}^{T}\!+\!\bar{\mathbf{I}}&\sum_{c}\mathbf{L}\bm{\vec{s}}^{(c)}\\ \sum_{c}(\bm{\vec{s}}^{(c)})^{T}\mathbf{L}^{T}&C\end{pmatrix}.$ (87)